OpenStudy (anonymous):

How to simplify this expression : sqrt(x^2-4)*sqrt(x^2+4) ?

OpenStudy (anonymous):

=sqrt(x^4 -16)

OpenStudy (anonymous):

\[ \sqrt{x-2} \times \sqrt{x+2} \times \sqrt{x+2} \times \sqrt{2}\]\[\sqrt{x-2} \times (\sqrt{x+2})^2 \times \sqrt{2}\]\[x+2 \times \sqrt{(x-2)2}~~~~~~~->~~~~~~x+2 \times \sqrt{x^2-4}\]

OpenStudy (anonymous):

I took out of the square root as much as I can. xiem2xoay just factored it.

OpenStudy (tkhunny):

@fbi2015 No, you took out more than you could. Please reconsider.

OpenStudy (anonymous):

\[\sqrt{x^2-4} \times \sqrt{x^2+4}=\sqrt{x^2-16}~~~~~~I~~~see~~~why.\]